OpenStudy (anonymous):

prove the following by mathmatical induction: 3+7+11+...+(4-1)=n(2n+1)

OpenStudy (jhannybean):

4=n(2n+1) ?

OpenStudy (anonymous):

oops, the left side is (4n-1)

OpenStudy (jhannybean):

\[\large 4n-1=n(2n+1)\] solving for n?

OpenStudy (jhannybean):

Oh wait nvm.... not solving for n.

OpenStudy (jhannybean):

Yeah.... how did you solve that now...lol

OpenStudy (dan815):

any ideas lol

OpenStudy (jhannybean):

oh i see, because n = n+1 you could easily substitute the "in for 4(n=1) which is basically an and an - 1 = 4(n+1)-1 so it's saying an + an-1 = 2an...... i think.

OpenStudy (dan815):

im giving up i need to look up what induction is!

OpenStudy (jhannybean):

ahh...

OpenStudy (dan815):

oh HEY! i think i got it!

OpenStudy (jhannybean):

yeah?

OpenStudy (dan815):

LHS: RHS let S be the sequence for n = 1 4(1)-1=3 : 1*2*1+3 =3 s(n+1)=(n+1)(2*(n+1)+1 =2n^2+5n+3 =2n^2+n+4n+3 =n(2n+1)+(4n+3) =S(n)+4(n+1)-1.... done

OpenStudy (dan815):

does that make sense?

OpenStudy (jhannybean):

Yeah... thats how i went about solving it but i didn't want it to be confusing with how you were explaining it. i got the 2n^2+5n+3

OpenStudy (dan815):

oh i dunno how to do induction lol i just looked up its meaning

OpenStudy (jhannybean):

Its kind of recursive in the sense that there's always that (4n+1) there somewhere in the problem.

OpenStudy (dan815):

show me the right formal for induction i need to see

OpenStudy (dan815):

format*

OpenStudy (anonymous):

Yes the idea is if its true for the base case and you can show that if its true for n then its true for n + 1. You can reach any case by applying the implication starting from the base case over and over until you reach the given index.

OpenStudy (jhannybean):

Whoah, too complicated. :|

OpenStudy (zzr0ck3r):

show its true for 1 assume its true for some n show its true(with that assumption) for n+1

OpenStudy (zzr0ck3r):

I can write it out and scan it but it might take me a min, have you solved it yet? I need to use sigma notation and that would take me to long on here.

OpenStudy (jhannybean):

scan it? lmao xD

OpenStudy (dan815):

im pretty sure my solution works, but i dont think its proper format

OpenStudy (anonymous):

This is how induction works. Why does induction give us that a given property is true for all \(n\). Lets say \(P(n)\) is a predicate which is true if a given property holds for \(n\). So if we can show that \(P(n)\) is true for \(P(1)\) and that \(P(n) \implies P(n + 1)\) then we can reach any given \(n\) by recursively applying the implication starting from the base case \(1\). So why does it hold for 3 just as an example? Because \(P(1) \implies P(2)\) and therefore \(P(2) \implies P(3)\).

OpenStudy (anonymous):

Yes and for the original poster, if you carefully read my discussion on induction you will understand why these proofs make sense to begin with.

OpenStudy (zzr0ck3r):

its dominoes, if one falls and then the next falls, then the rest will fall.

OpenStudy (anonymous):

Actually that's a nice visual description of it.

OpenStudy (zzr0ck3r):

@comput313 let me know if you have any questions about the proof.

OpenStudy (zzr0ck3r):

ok it is true for n =1 assume sigma from 1 to n (4i-1) = n(2n+1) then sigma from 1 to n+1 (4i-1) = 4(n+1) + sigma from 1 to n of (4i-1) = 4(n+1) -1+n(2n+1) = 2n^2+5n+3 = (n+1)(2(n+1) + 1) thus by induction it is true for all n

OpenStudy (zzr0ck3r):

sorry typo fixed

OpenStudy (anonymous):

You can use latex here =P. $$\displaystyle\large\sum_{i=1}^n$$

OpenStudy (zzr0ck3r):

I know the math not the latek:)

OpenStudy (zzr0ck3r):

\[\sum_{i = 1}^{n} ( 4i-1) and \sum_{i=1}^{n+1} (4i-1)\]

OpenStudy (jhannybean):

@zzr0ck3r rewriting.\[T \to n=1\]\[\large \sum_{n=1}^{n}(4n-1) = n(2n+1)\]\

OpenStudy (jhannybean):

Haha aww.... ok. I shall let you continue. LaTex is fun.

OpenStudy (anonymous):

Indeed, I used LaTeX for all of my problem set solutions for a course this year.

OpenStudy (zzr0ck3r):

\[\sum_{i =1}^{n} (4i-1) = n(2n+1)\] then \[\sum_{i=1}^{n+1}(4i-1) = 4(n+1)-1+\sum_{i=1}^{n}(4i-1) = \] 4(n+1) -1+n(2n+1) = 2n^2+5n+3 = (n+1)(2(n+1) + 1) qed.

OpenStudy (zzr0ck3r):

@comput313

OpenStudy (anonymous):

No, he's probably long gone.

OpenStudy (zzr0ck3r):

bah

OpenStudy (dan815):

oh thanks zzrocker :)

OpenStudy (dan815):

now i understand induction

OpenStudy (jhannybean):

\[\large \sum_{n=1}^{n}(4n-1) = n(2n+1)\]\[\large \sum_{n=1}^{n+1}(4n-1)=4(n+1)+\sum_{n=1}^{n}(4n-1)\]\[\large = 4(n+1) -1+n(2n+1) = 2n^2+5n+3 = (n+1)(2(n+1) + 1)\]

OpenStudy (jhannybean):

mwuahahaha

OpenStudy (zzr0ck3r):

good deal. you will need it for any 300+ level math class.

OpenStudy (jhannybean):

Super Saiyan :P Jk..

OpenStudy (dan815):

well iam only 12 years old, so i have time

OpenStudy (dan815):

12 in elephant years

OpenStudy (zzr0ck3r):

let n = 1 then for sure 3 = 3 so assume its true for some n thus \[\sum_{i=1}^{n}(4i-1) = n(2n+1)\] consider n+1 then \[\sum_{i=1}^{n+1}(4i-1) = 4(n+1)-1+\sum_{i=1}^{n}(4i-1)\] we know from out assumption about n =1 that this is equal to \[4(n+1)-1+n(2n+1) = (n+1)(2(n+1)+1)\] thus by induction we have shown that if it is true for n it is true for n+1 and since its true for n =1 it follows that it is true for all n in N.

OpenStudy (zzr0ck3r):

yes lots of time.:)

OpenStudy (jhannybean):

omg i forgot the "-1" after "4(n-1)" -_____- oh mygoodness.